OpenStudy (anonymous):

Solve: 0 = 4x^2 - 8x - 5?

OpenStudy (callisto):

0 = 4x^2 - 8x - 5 0 = (2x+1)(2x-5) (2x+1) =0 or (2x-5)=0 Can you solve it from here?

OpenStudy (anonymous):

not really can you walk me through it?

OpenStudy (callisto):

Okay, first, do you know how to do factorization?

OpenStudy (anonymous):

sort of

OpenStudy (anonymous):

I know Greatest common factors but never really learned group factoring or poly/trionomial factoring

OpenStudy (callisto):

Oh... have you learnt quadratic formula?

OpenStudy (anonymous):

nope

OpenStudy (callisto):

Then....Normally, how do you solve quadratic equation?

OpenStudy (anonymous):

Guesstimate

OpenStudy (callisto):

Do you know how to do completing square?

OpenStudy (anonymous):

dont you turn ax2 + bx + c = 0 to a(x+d)2 + e = 0

OpenStudy (callisto):

Yes

OpenStudy (anonymous):

I know the formats but not really how to convert it

OpenStudy (callisto):

So, let's use that method, shall we start?

OpenStudy (anonymous):

sure

OpenStudy (callisto):

0 = 4x^2 - 8x - 5 5 = 4x^2 - 8x Understand?

OpenStudy (anonymous):

I thought it would be -5 not just 5?

OpenStudy (callisto):

Nope add 5 on both sides to the equation 0 = 4x^2 - 8x - 5. 0+5 = 4x^2 - 8x - 5 +5 5 = 4x^2 - 8x

OpenStudy (anonymous):

alright got it

OpenStudy (anonymous):

then what?

OpenStudy (callisto):

Take out common factor on the right side 5 = 4x^2 - 8x 5 = 4(x^2 - 2x) Got it?

OpenStudy (anonymous):

yea 2x is the common factor of 4 and 8 right?

OpenStudy (callisto):

Hmm.. Yes... But we need to complete square, so, keep it as x^2

OpenStudy (anonymous):

alright

OpenStudy (callisto):

Since a^2-2ab+b^2 = (a-b)^2 now, a=x So, we have x^2 -2xb + b^2 Compare this to the equation we have ( Consider the things inside the bracket first) To complete square: (x^2 - 2x) = x^2 -2x + (2/2)^2 - (2/2)^2 = x^2 -2x +1 -1 = (x-1)^2 -1

OpenStudy (callisto):

Got it so far?

OpenStudy (anonymous):

lagged sorry

OpenStudy (callisto):

no worry. But I have to leave in 30 minutes. just a notice. Tell me if you understand this step or not after you've finished

OpenStudy (anonymous):

I don't completely understand this step

OpenStudy (callisto):

So, which one you don't know?

OpenStudy (anonymous):

first 1

OpenStudy (callisto):

Can you copy that step?

OpenStudy (anonymous):

Since a^2-2ab+b^2 = (a-b)^2 now, a=x So, we have x^2 -2xb + b^2

OpenStudy (callisto):

a^2-2ab+b^2 = (a-b)^2 <- this is the identity for perfect square..Got it?

OpenStudy (anonymous):

yea

OpenStudy (callisto):

Now compare the two, what do you observe? a^2-2ab+b^2 (x^2 - 2x)

OpenStudy (anonymous):

that ^2 is in both

OpenStudy (callisto):

So, we can say x ~ a , if we want to get the perfect square, what is the value of b required?

OpenStudy (callisto):

Hints: compare -2ab and -2x, take x=a as I've mentioned

OpenStudy (anonymous):

-2x is common factor?

OpenStudy (callisto):

Yes, you can say so

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then what?

OpenStudy (callisto):

What is the value of b you get?

OpenStudy (anonymous):

-2?

OpenStudy (callisto):

Nope... -2ab = -2x Put a=x -2xb= -2x b=1 Got it?

OpenStudy (anonymous):

yea

OpenStudy (callisto):

So... (x^2 - 2x) = x^2 - 2x +1 -1 Got it?

OpenStudy (anonymous):

yep

OpenStudy (callisto):

(x^2 - 2x) = x^2 - 2x +1 -1 = (x-1)^2 -1 Got it?

OpenStudy (callisto):

Since (x-1)^2 = x^2 -2x +1

OpenStudy (callisto):

@Brad1996 ?

OpenStudy (anonymous):

im back

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

yea i got it

OpenStudy (anonymous):

1hr on 1 question :l

OpenStudy (callisto):

Back to the previous steps, we have 5 = 4x^2 - 8x 5 = 4(x^2 - 2x) Put (x^2 - 2x) =(x-1)^2 -1 into 5 = 4(x^2 - 2x) We'll get 5 = 4 [(x-1)^2 -1] Got it so far?

OpenStudy (callisto):

lol nearly finish :)

OpenStudy (anonymous):

yep

OpenStudy (callisto):

Divide both sides by 4 5/4= 4 [(x-1)^2 -1] /4 5/4 = [(x-1)^2 -1] Got it?

OpenStudy (swag):

x = -1/2, 5/2

OpenStudy (callisto):

@SWAG Answer is not that important. The asker needs to know how to do it, instead of knowing the numbers for a single question

OpenStudy (swag):

You showed him how to do it though

OpenStudy (callisto):

Haven't reached the last step. There's still possibility that the asker doesn't understand

OpenStudy (callisto):

@Brad1996 Do you understand the previous steps ? Divide both sides by 4 5/4= 4 [(x-1)^2 -1] /4 5/4 = [(x-1)^2 -1]

OpenStudy (anonymous):

yea

OpenStudy (callisto):

Add 1 to both sides. 5/4 +1= (x-1)^2 -1 +1 9/4 = (x-1)^2 <- Got it?

OpenStudy (anonymous):

yep

OpenStudy (callisto):

Now take square root for both sides. 9/4 = (x-1)^2 x-1 = 3/2 or x-1 = -3/2 Note that \[\sqrt{y^2} = \pm y\] Got it?

OpenStudy (anonymous):

yea

OpenStudy (callisto):

Solve the 2 equations and you'll find x :) can you do it?

OpenStudy (anonymous):

yep

OpenStudy (callisto):

What do you get for the answers?